B.      Calculate the payback period for a project that requires an investment of $120,000, and has the following returns over a 4 year period:  
Year 1 2 3 4
Cash Flow $32,456 $44,652 $39,673 $38,950
C.      Calculate the payback period for the project with the following cash flow projection. What do you expect might happen if the time value of money is applied to this project?
Year 0 1 2 3 4
CF ($235,498) $64,987 $54,378 $49,987 $43,675
D.     Select the preferable project using the payback period method. Under what conditions does the use of the payback period make sense for project selection?
0 1 2 3 4
CF Project 1 ($98,024) $43,423 $35,674 $23,456 $18,316
CF Project 2 ($79,315) $26,000 $26,000 $26,000 $26,000
Answer #1
Payback period
B.Investment 120000
Year Cashflow Cumulative Cashflow
0 -120000 -120000
1 32456 -87544
2 44652 -42892
3 39673 -3219
4 38950
Payback Period = 3 years + 3219/38950 = 3.08 years
c.Investment 120000
Year Cashflow Cumulative Cashflow
0 -235498 -235498
1 64987 -170511
2 54378 -116133
3 49987 -66146
4 43675 -22471
Payback Period = More than 3 years if there is cashflow
after that else no payback period
Time Value means discounted payback period .If the
cashflows are discounted it will take further more
period to Payback.
c.Project 1 & 2
Project 1
Year Cashflow Cumulative Cashflow
0 -98024 -98024
1 43423 -54601
2 35674 -18927
3 23456 4529
4 18316
Payback Period = 3 years + 4529/18316 = 3.25 years
Project 2
Year Cashflow Cumulative Cashflow
0 -79315 -79315
1 26000 -53315
2 26000 -27315
3 26000 -1315
4 26000
Payback Period = 3 years + 1315/26000 = 3.05 years
Payback period is generally used when cashbudget
is more important to choose a project.The one
which returns capital earlier would be chosen.